OpenStudy (anonymous):

The credit remaining on a phone card (in dollars) is a linear function of the total calling time made with the card (in minutes). The remaining credit after 35 minutes of calls is $24.40, and the remaining credit after 71 minutes of calls is $18.64 . What is the remaining credit after 104 minutes of calls?

OpenStudy (zpupster):

are you there??

OpenStudy (anonymous):

yeah im here! :D

OpenStudy (zpupster):

this is a little involved since it is a line we need the slope look at the question and come up with 2 ordered pairs (x,y)

OpenStudy (anonymous):

how about (7,5)

OpenStudy (anonymous):

oh were you asking me to make one ?

OpenStudy (anonymous):

)Point A is located at negative 7 over 10 and point B is located at negative 2 over 10. What is the distance between points A and B?

OpenStudy (anonymous):

what sisi?

OpenStudy (zpupster):

no from the question

OpenStudy (zpupster):

hhow about (35,24.40) and (71,18.64) our third would be (104, y)

OpenStudy (zpupster):

\[m=\frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

oh okidoke !

OpenStudy (zpupster):

we will use the above formula to find a slope can oyu plug the values in??

OpenStudy (anonymous):

would i basically just use the last two points?

OpenStudy (zpupster):

x2=35 x1=71 and y2=24.40 y1 = 18.64 do oyu see now??

OpenStudy (anonymous):

yeuup! m= y1-y2/x1-x2 m=18.64-24.40/71-35

OpenStudy (anonymous):

m=-5.76/36 now what?

OpenStudy (zpupster):

ok do the math

OpenStudy (anonymous):

-.16

OpenStudy (anonymous):

thats what i got at least

OpenStudy (zpupster):

m=-5.76/36=-.16 ok?

OpenStudy (anonymous):

yeup!

OpenStudy (zpupster):

now we need to use the point slope form \[y-y1 = m(x-x1)\]

OpenStudy (zpupster):

use (35,24.40) \[y-24.4=-.16(x-35)\]

OpenStudy (zpupster):

you nedd to put it in this slope intercept form form \[y=mx+b\]

OpenStudy (anonymous):

sorry i'm doing two things at once here okidoke im not quite sure where to go from here im at y-24.4=-.16x+5.6

OpenStudy (anonymous):

because i did the math on the one side

OpenStudy (zpupster):

add 24.4 to both sides

OpenStudy (anonymous):

oh right ah ! okidoke y=-.16x=30

OpenStudy (zpupster):

now we have our line and we know x = 104 for our third point so plug it in the y=-.16(104) +30 what do you get??

OpenStudy (anonymous):

46.64!

OpenStudy (anonymous):

but that doesn't make any sensee

OpenStudy (zpupster):

try one more time

OpenStudy (anonymous):

oh dangit okidoke uhh 13.36!!

OpenStudy (zpupster):

yep and that is your answer good job!! you can see the point falls on our line-graph

OpenStudy (anonymous):

awesome thanks so much dude ! :DD

OpenStudy (zpupster):

no problem!! i would advise you glance over this again and pay attention to the steps.

OpenStudy (anonymous):

For sure thanks again! :)